A ball is thrown horizontally at 8m/s. (a) how fast is it moving 2second later? In what direction is it moving?
1
Expert's answer
2021-10-15T10:22:49-0400

(a) The horizontal component of the speed remains the same during the entire flight, the vertical component:

"v_y=gt=9.8·2=19.6\text{ m/s},\\
v=\sqrt{v_y^2+v_x^2}=21.2\text{ m/s}."


The direction is


"\theta=\arctan\frac{v_y}{v_x}=67.8°"

below the horizontal.